OpenStudy (faf3):

a 2kg blob of putty moving at 3 m/s slams into a 2 kg of putty at rest. what is the initial momentum of the two blobs? what must be the final momentum of the two blobs? Also, calculate the velocity of the two stuck together blobs of putty after the collision.

OpenStudy (anonymous):

Use the conservation of momentum equation m1v1+m2v2=(m1+m2)Vf to solve for Vf. Then momentum p is just p=(m1+m2)Vf .

OpenStudy (anonymous):

6